OpenStudy (anonymous):

HELP!!!!! What is the vertex in the quadratic function y = - 2x^2+ 12x - 4 ? Write your answer as an ordered pair.

OpenStudy (zzr0ck3r):

x coordinate of the vertex is given by -b/(2a) what is a and what is b?

OpenStudy (anonymous):

A Is 2x^2 and b is 12 ???

OpenStudy (zzr0ck3r):

a = -2 b = 12 we just take the number in fron of the x

OpenStudy (zzr0ck3r):

so \[\frac{-12}{2(-2)} = ?\]

OpenStudy (anonymous):

ohh okay.. and is it 3

OpenStudy (anonymous):

??

Directrix (directrix):

@xoxox123 Yes, the x-coordinate of the parabola is 3. Now, substitute x = 3 into y = - 2x^2+ 12x - 4 to get the coordinate of the y value of the vertex.

OpenStudy (anonymous):

(3,50) I guess

OpenStudy (anonymous):

is y -4?

OpenStudy (anonymous):

oh sorry i was mistaken the answer is (3,14)

Directrix (directrix):

@killua_vongoladecimo Code of conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

can someone help me find y ?

OpenStudy (anonymous):

i'm sorry.....well then here is the solution To get the x-coordinate of the vertex use the equation \[\frac{ -b }{ 2a }\] so it would be\[\frac{ -12 }{ -4 }\] or 3 so it is the x-coordinate. To get the y-coordinate equate y to \[y=-2x^{2} + 12x - 4\] Then substitute 3 to x which results to -18 + 36 - 4 or 14 So the coordinates of the vertex are (3,14)

OpenStudy (anonymous):

how would you do this one ?? What is the vertex in the quadratic function y = - 4x^2 + 8x - 3 ? Write your answer as an ordered pair.

OpenStudy (anonymous):

@killua_vongoladecimo